The Elements of Theses

by Perry H. Disdainful

B.A. (University of Northern South Dakota at Hoople) 1978 M.S. (Ed’s School of Quantum Mechanics and Muffler Repair Shop) 1989

A dissertation submitted in partial satisfaction of the

requirements for the degree of

Doctor of Philosophy

in Aboriginal Basketry

in the

GRADUATE DIVISION

of the

UNIVERSITY of CALIFORNIA at BERKELEY

Committee in charge:

Professor Ignatius Arrogant, Chair Professor Ivory Insular Professor General Reference

The dissertation of Perry H. Disdainful is approved:

Chair Date \overline{Date} \overline{Date}

University of California at Berkeley The Elements of Theses

Copyright 1994 by Perry H. Disdainful

Abstract

The Elements of Theses

by Perry H. Disdainful

Doctor of Philosophy in Aboriginal Basketry University of California at Berkeley Professor Ignatius Arrogant, Chair

Theses have elements. Isn’t that nice?

Professor Ignatius Arrogant Dissertation Committee Chair

To myself, Perry H. Disdainful,

the only person worthy of my company.

Contents

List of Figures v List of Tables vi

I First Part 1

1 Introduction 2 2 Previous Work 4 3 Conclusion 6 Bibliography 7 A Some Ancillary Stuff 9

List of Figures

2.1 A first figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 A second figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

List of Tables

1.1 A normalsize table. There has been a complaint that table captions are not

single-spaced. This is odd because the code indicates that they should be. . 2 1.2 A small table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Acknowledgements

I want to “thank” my committee, without whose ridiculous demands, I would have graduated so, so, very much faster.

1

Part I First Part

Chapter 1

Introduction

Every dissertation should have an introduction. You might not realize it, but the

introduction should introduce the concepts, backgrouand, and goals of the dissertation.

Title  Author
War And Peace  Leo Tolstoy
The Great Gatsby F. Scott Fitzgerald

Table 1.1: A normalsize table. There has been a complaint that table captions are not single-spaced. This is odd because the code indicates that they should be.

Table 1.2: A small table.

Title  Author
War And Peace  Leo Tolstoy
The Great Gatsby F. Scott Fitzgerald

Chapter 2

Previous Work

Some other research was once performed.

Figure 2.1: A first figure.

5

Figure 2.2: A second figure.

6

Chapter 3 Conclusion Bibliography

[1] Jacques Désarménien. How to run TEX in french. Technical Report SATN-CS-1013, Computer Science Department, Stanford University, Stanford, California, August 1984.

[2] David Fuchs. The format of TEX’s DVI files version 1. TUGboat, 2(2):12–16, July 1981.

[3] David Fuchs. Device independent file format. TUGboat, 3(2):14–19, October 1982.

[4] Richard K. Furuta and Pierre A. MacKay. Two TEX implementations for the IBM PC. Dr. Dobb’s Journal, 10(9):80–91, September 1985.

[5] Donald E. Knuth. The WEB system for structured documentation, version 2.3. Technical Report STAN-CS-83-980, Computer Science Department, Stanford University, Stanford, California, September 1983.

[6] Donald E. Knuth. The TEX Book. Addison-Wesley, Reading, Massachusetts, 1984. Reprinted as Vol. A of Computers & Typesetting, 1986.

[7] Donald E. Knuth. Literate programming. The Computer Journal, 27(2):97–111, May [8] Donald E. Knuth. A torture test for TEX, version 1.3. Technical Report STAN-

CS-84-1027, Computer Science Department, Stanford University, Stanford, California, November 1984.

[9] Donald E. Knuth. TEX: The Program, volume B of Computers & Typesetting. Addison-Wesley, Reading, Massachusetts, 1986.

[10] Leslie Lamport. L^{A}TEX: A Document Preparation System. User’s Guide and Reference Manual. Addison-Wesley, Reading, Massachusetts, 1986.

[11] Oren Patashnik. BibTEXing. Computer Science Department, Stanford University, Stanford, California, January 1988. Available in the BibTEX release.

[12] Oren Patashnik. Designing BibTEX Styles. Computer Science Department, Stanford University, January 1988.

[13] Arthur L. Samuel. First grade TEX: A beginner’s TEX manual. Technical Report SATN-CS-83-985, Computer Science Department, Stanford University, Stanford, California, November 1983.

[14] Michael D. Spivak. The Joy of TEX. American Mathematical Society, 1985.

Appendix A

Some Ancillary Stuff

Ancillary material should be put in appendices, which appear after the bibliogra-

phy.